Grade 12th passAlgebra

if x^a=y^b=z^c , then find the value of b. and it is given that y^2=xz.

rohan singh
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
Dear Rohan
 
y^2 = xz
 
k^2/b = k^1/a * k^1/c
 
hnce
 
2/b = 1/a + 1/c
 
Hence
 
b = 2ac / (a + c)
 
Regards
Arun (askIITians forumexpert)